### v2

# MultiStream-LLM: Bridging Modalities for Robust Sign Language Translation

Abstract
Despite progress in gloss-free Sign Language Translation (SLT), monolithic end-to-end models consistently fail on two critical components of natural signing: the precise recognition of high-speed fingerspelling and the integration of asynchronous non-manual cues from the face. Recent progress in Automated Sign Language Translation with Large Language Models has side stepped this challenge, forcing a single network to learn these simultaneously resulting in poor performance when tasked with translating crucial information such as names,places, and technical terms.
We introduce MultiStream-LLM, a modular framework designed to overcome these limitations. Our approach employs separate, specialized predictors for continuous signing, fingerspelling, and lipreading. Each expert network first decodes its specific modality into a sequence of tokens. These parallel streams are then fused by a lightweight transformer that resolves temporal misalignments before passing the combined representation to a Large Language Model (LLM) for final sentence generation.
Our method establishes a new state-of-the-art on the How2Sign benchmark with a BLEU-4 score of 23.5 and achieves 73.2% letter accuracy on the challenging ChicagoFSWildPlus fingerspelling dataset. These results validate our core hypothesis: by isolating and solving distinct recognition tasks before fusion, our multi-expert approach provides a more powerful and effective pathway to robust, high-fidelity sign language translation.
Introduction
Sign languages are fully fledged natural languages that are used by millions of Deaf and hard-of-hearing people around the world to communicate. They make use of both manual gestures (handshapes, movements and locations) alongside non‐manual signals (facial expressions, mouthings, and mouth gestures [^1] In sign language linguistics, mouthing denotes silent articulation of spoken words that accompany signs, while mouth gestures are nonverbal movements integral to the sign itself, conveying grammatical or affective nuance). They vary from country to country, for example: American Sign Language (ASL), British Sign Language (BSL) and German Sign Language (DGS) (HandSpeak 2024; British Sign Language 2024; GebärdenVerstehen 2024) where each possesses its own lexicon and grammar, even if many share spatial and syntactic parallels.
Early methods focused on Isolated Sign Language Recognition (ISLR) but relied on expertly annotated datasets. These annotations include “glosses (Camgoz et al. 2018)” which are English translations for individual words. More recent works have demonstrated that glosses are no longer a strict requirement for effective SLT with the introduction of large, gloss-free datasets (e.g. YouTube‐ASL (Uthus, Tanzer, and Georg 2023), Public DGS (Hanke et al. 2020), How2Sign (Duarte et al. 2021)) as well as increased model size and the use of LLM’s (Lin et al. 2023). However even with ChicagoFSWildPlus (B. Shi and Livescu 2019) a dataset focused on ASL fingerspelling modern approaches often struggle with two key challenges:
  • Fingerspelling: Used for proper names, technical terms and loanwords’ fingerspelling makes up roughly 12–35% of ASL, 8–20% of BSL, and 5–10% of DGS (Valli et al. 2011; Pfau, Steinbach, and Woll 2012). Despite its importance, it remains under‐supported in gloss‐free SLT systems like Llava-SLT (Liang et al. 2024) where it is often treated as just another sub‐task of the full translation pipeline, which leads to:
    • Temporal blurring. Consecutive letters are articulated rapidly, with hand‐shape transitions that blur together; end‐to‐end models struggle to segment the stream into discrete letters without explicit supervision.
    • Visual ambiguity. Many handshapes differ only subtly (e.g., “M” vs. “N”), and without a dedicated fingerspelling detector these fine‐grained distinctions get lost in the broader sign‐recognition feature space.
    • Data imbalance. Fingerspelling accounts for only a fraction of all signing in available corpora, so a single network trained on full sentence translation tends to under‐fit the less frequent letter sequences.
    • As a result, fingerspelled terms in gloss-free SLT are frequently mistranslated or dropped entirely.
  • Non‐manual cues Mouthings and facial expressions carry critical disambiguating signals, akin to phonemes or visemes in speech (Encyclopædia Britannica 2024; Ozel 2024) but existing SLT models rarely exploit them, mostly due to the asynchronous nature of these signals.
In parallel, breakthroughs in Automatic Speech Recognition (ASR) (Long et al. 2025) and Visual Speech Recognition (VSR, i.e., lipreading) (Thomas, Fish, and Bowden 2025) demonstrate how large scale pretraining and powerful architectures such as ViTs (Dosovitskiy et al. 2021) and LLM decoders (Radford et al. 2022) can deliver human‐level performance on spoken language tasks. Yet sign language systems lag behind, largely because (a) large, freely available signed corpora are scarce; (b) the spatiotemporal complexity of co‐articulated gestures exceeds that of spoken‐language signals; and (c) unlike speech, signing and mouthing are often offset in time especially during fingerspelling, where manual letters may precede or follow their mouthings.
To address these gaps, we introduce MultiStream‐LLM, a gloss‐free SLT framework that:
  1. 1.
    Integrates three modalities continuous sign recognition, fingerspelling detection, and lipreading.
  2. 2.
    Fuses their complementary signals via a lightweight transformer encoder that explicitly handles temporal asynchronicity.
  3. 3.
    Decodes the fused representation with a fine‐tuned Large Language Model, yielding contextually accurate translations without any direct gloss supervision.
We evaluate MultiStream‐LLM on YouTube‐ASL and ChicagoFSWildPlus, achieving state‐of‐the‐art results: sequence classification accuracy of 99.2%, letter accuracy of 73.2%, and BLEU‐4 of 23.5. Our approach demonstrates that (1) non‐manual cues especially lipreading—significantly boost gloss‐free SLT, and (2) combining modality‐specific predictors with an LLM yields scalable, high‐quality translation for real‐world signed content.
Related Work
Sign Language Understanding
Sign Language Understanding (SLU) has evolved through three increasingly ambitious tasks:
  1. 1.
    Isolated Sign Language Recognition (ISLR): Early work focused on recognizing individual signs in controlled settings. Methods such as 3D CNNs over spatiotemporal volumes (Min et al. 2021) and hybrid CNN–HMM pipelines (Koller et al. 2019) achieved high accuracy on small vocabularies (e.g., 100–500 signs), but relied on carefully segmented clips and extensive manual gloss annotations.
  2. 2.
    Continuous Sign Language Recognition (CSLR): To scale beyond isolated signs, CSLR methods adopted sequence‐level training with Connectionist Temporal Classification (CTC) loss (Graves et al. 2006). Recurrent architectures (CNN–BiLSTM–CTC) (Graves and Schmidhuber 2005) and, more recently, Transformer encoders (Min et al. 2021) enabled end‐to‐end gloss prediction from unsegmented video. While these models tolerate variable sign durations, they still depend on gloss‐level transcripts and struggle with out‐of‐vocabulary signs and fingerspelling.
  3. 3.
    Sign Language Translation (SLT): SLT goes one step further by generating natural language sentences directly from sign video.
    • Gloss‐based SLT pipelines first predict a gloss sequence using CSLR and then translate glosses to text via a sequence‐to‐sequence model (Camgoz et al. 2018). Examples include CNN–LSTM–attention frameworks that achieve fluent translation but inherit errors from the gloss recognizer.
    • Gloss‐free SLT bypasses intermediate glosses, learning a direct video‐to‐text mapping. Recent works leverage large‐scale video–text pretraining (Zhou et al. 2023; Lin et al. 2023; Zhou et al. 2024; Tarrés et al. 2023) and multimodal contrastive objectives (Gueuwou et al. 2025) to improve translation fluency. More recent approaches have focussed on leveraging LLM’s language modelling capabilities to incorporate sign features either from RGB (uthus2024youtubeasl), Pose (Fish and Bowden 2025) , or a fusion of both modalities (Li et al. 2025; Wong, Camgoz, and Bowden 2024). However, these models often underperform on fingerspelling and non‐manual cues, since they attempt to learn all visual subtasks in a single end‐to‐end network.
Lipreading and Visual Speech Recognition
Lipreading, or visual speech recognition (VSR), can complement sign language understanding by capturing silent speech cues. Modern systems employ deep learning-based architectures. LipNet (Assael et al. 2016) introduced an end-to-end model using spatiotemporal CNNs (Lecun et al. 1998; Carreira and Zisserman 2018) and BiLSTMs (Graves and Schmidhuber 2005), achieving state-of-the-art sentence-level lipreading performance. Transformer-based models, such as Lipformer (Xue et al. 2023), further improved contextual understanding across video frames.
Recent advances leverage self-supervised learning and multimodal fusion to enhance lipreading robustness. AV-HuBERT (Shi et al. 2022b) demonstrated the efficacy of large-scale audiovisual pretraining, achieving strong performance under challenging conditions (Thomas, Fish, and Bowden 2025). However, lipreading remains underexplored in the context of fingerspelling and full‐sentence SLT, dispite mouthing cues carrying up to 40% of linguistic information in natural sign production (e.g., lexical disambiguation, grammatical markers) (Aparicio et al. 2017).
Figure 1: An overview of our approach. First we extract facial regions from 16 frames of input video. We apply random pixel masking at 50% probability, add positional embedding, and encode visual features via a vision transformer encoder (ViT-base (Dosovitskiy et al. 2021)). We implement temporal downsampling via 1D convolution and then use a Connectionist Temporal Classification (CTC) linear head to predict sequences of phonemes. During training, we also fine-tune an LLM (Vaswani et al. 2023) (Llama (Touvron et al. 2023)) to reconstruct words from phonemes with a text-only dataset (Carnegie Mellon University 2014). During inference, the phonemes from the CTC (Graves et al. 2006) head are processed via the LLM to reconstruct the predicted text. This can be performed end-to-end or in two stages, depending on available resources.
Multimodal Approaches
Multimodal learning has emerged as a promising approach to improve fingerspelling detection by integrating visual, linguistic, and contextual information. Previous systems fused RGB video, pose sequences, and optical flow to enhance recognition accuracy. Recent works such as GFSLT-VLP (Zhou et al. 2023) incorporated video-text contrastive learning and cross-modal attention mechanisms to bridge the gap between sign language and natural language representations.
However, current fingerspelling detection methods remain focused on hand centric cues and fail to leverage the rich information present in lip movements. Approaches such as synthetic data augmentation (Fowley and Ventresque 2022), large “in‐the‐wild” collections (Zhidebayeva et al. 2024; Georg et al. 2024), and pose‐based models achieve moderate accuracy on isolated letter streams but do not integrate mouthings. This design overlooks the well‐documented co‐articulation between handshapes and lip movements in natural signing (Aparicio et al. 2017), leading to persistent errors on homonymous handshapes (e.g., “M” vs. “N”) and proper‐noun detection. To date, no SLT system has explicitly fused high‐resolution fingerspelling and lipreading streams while accounting for their temporal misalignment.
LLMs in SLT and VSR
The rise of Large Language Models (LLMs) has revolutionized sign language understanding, fingerspelling detection, and lipreading by leveraging vast linguistic knowledge and multimodal learning capabilities. In SLT, LLMs have been employed to bridge the gap between sign video sequences and natural language text. Newer approaches (Ruiz and Martinez 2025) applied transformer-based architectures to generate coherent text from sign videos, outperforming traditional gloss-based approaches. GFSLT-VLP (Zhou et al. 2023) further enhanced gloss-free SLT performance.
For fingerspelling detection, LLMs have facilitated cross-modal learning by aligning hand gestures with corresponding textual representations (Tanzer 2024; Shi 2023). These approaches demonstrated that LLMs can effectively enhance fingerspelling recognition by integrating linguistic and visual modalities.
In lipreading, LLMs have been integrated into audiovisual speech recognition pipelines to improve contextual understanding and transcription accuracy. The first approach to utilize LLM-based encoders (Yeo et al. 2024) enhanced lipreading robustness, even under challenging conditions. These advancements underscore the potential of LLMs to bridge modality gaps and improve performance across SLT, fingerspelling, and lipreading tasks.
Method
Our proposed framework integrates lipreading, fingerspelling detection, and sign language recognition to achieve robust sign language translation. The architecture as shown in Figure 1, comprises four stages: (1) Sequence Type Classification, (2) Modality-Specific Predictors, (3) Multimodal Fusion and (4) Final Prediction.
$f:\mathbf{X}=\{x_{1},\dots,x_{T}\}$ (1)
[…]
$\xrightarrow{\text{Fusion}}\quad\mathbf{S}=\{s_{1},\dots,s_{K}\}$
Equation 1 shows that given a video input $X$ we aim to learn a number of functions for the individual tasks described in the introduction. $L_{fs}$, $L_{sign}$ and $L_{lip}$ decode finger spellings, signing and lip reading respectively. These decoding’s are then fused together and passed through an LLM to get the final translation.
Pre-Processing Pipeline
In the pre-processing stage, the input videos are segmented into two distinct video sequences using MediaPipe (Lugaresi et al. 2019). The first sequence consists of cropped frames that contain just the hands, $X_{hands}$. The second sequence consists of cropped frames that contain just the face, $X_{face}$. Both video sequences are resized to a consistent resolution of $224\times 224$ pixels per frame. Each video is normalized and converted to batches of shape $B\times T\times 3\times 224\times 224$, where $B$ is the batch size, $T$ is the number of frames, and the last three dimensions represent the RGB channels, height, and width giving:
$X\in\mathbb{R}^{B\times T\times 3\times 224\times 224}$ (2)
As well as pre-processing the video inputs we also pre-process the corresponding labels. This entails Pseudo-Glossing to remove unneeded text from labels to reduce the required lexicon of the Sign Language Prediction model.
Sequence Type Classification
We classify each input sequence $X$ into one of the following classes {Continuos Sign, Fingerspelling, or Rest} via a ViT‐based encoder:
$F=\frac{1}{T}\sum_{t=1}^{T}\mathrm{ViT}(X_{hands_{t}})\;\in\;\mathbb{R}^{d_{\mathrm{model}}}$ (3)
We then pass the features into a linear layer with drop-out to classify the sequence into one of the three classes.
$\hat{Y}=W\,\tilde{F}+b\;\in\;\mathbb{R}^{\mathrm{num\_classes}}$ (4)
Where $\hat{Y}$ is the vector of class logits produced by the sequence‐type classifier.
To learn which type of sequence is currently active in the video, we implement a hard Gumbel–Softmax with temperature $\tau$ and Gumbel noise $u$ which yields a differentiable, one‐hot selection between the three classes:
$g_{i}$ $=\frac{\exp\bigl{(}(\hat{y}_{i}+u_{i})/\tau\bigr{)}}{\sum_{j=1}^{N_{\mathrm{classes}}}\exp\bigl{(}(\hat{y}_{j}+u_{j})/\tau\bigr{)}},$ (5)
$\mathbf{g}$ $=[g_{\mathrm{sign}},\,g_{\mathrm{fs}},\,g_{\mathrm{rest}}]$
where $g_{i}$ are the branch probabilities, if $g_{rest}$ is selected a unique rest vector is returned
Modality-Specific Predictors
Based on the predicted sequence type, the system dynamically selects the appropriate network to process either the fingerspelling or continuos sign alongside the lip reading from the video input $X$. We define three additional ‘expert’ networks that can be selected to extract these features
Manual Feature Prediction:
We model both continuous sign and fingerspelling using a DinoV2‐based visual backbone followed by a CTC head. Concretely, each input clip $X\in\mathbb{R}^{B\times T\times C\times H\times W}$ is flattened to BTB\,T frames and passed through a pretrained DINOv2 backbone to produce frame features. We then add a learned positional embedding and apply dropout. Finally, a linear classifier projects $\tilde{F}$ to logits over the $N_{\mathrm{classes}}$ tokens:
$F$ $=\mathrm{DinoV2Backbone}(X_{\mathrm{hands}})$ $\in\mathbb{R}^{B\times T\times d_{\mathrm{model}}},$ (6)
$P$ $\in\mathbb{R}^{1\times T_{\max}\times d_{\mathrm{model}}},$
$\tilde{F}$ $=\mathrm{Dropout}\bigl{(}F+P_{:,1:T,:}\bigr{)}$ $\in\mathbb{R}^{B\times T\times d_{\mathrm{model}}},$
$L_{n}$ $=\tilde{F}W_{n}+b_{n}$ $\in\mathbb{R}^{B\times T\times N_{n}}$
where $P$ are the positional embeddings and $n=[fs,sign]$
Lip Reading Model:
Simultaneously, we use a Lip Reading model which is comprised of a masked feature extractor, an adapter network and a CTC head. Given a batch of face‐cropped video frames $X$ we first apply random patch masking (50% drop rate) and add learnable positional embeddings. These masked frames are then encoded by a ViT to produce per‐frame feature vectors $Z$. Next we pass $Z$ through an adapter network consisting of a cascade of 1D convolutional layers and a pooling layer to downsample the temporal dimension while retaining an intermediate feature dimension $D_{adapter}$. Finally, a CTC head maps these downsampled features to phonemes $L$.
$\begin{split}Z&=\mathrm{ViT}(X_{face})\;\in\;\mathbb{R}^{B\times T\times D_{\mathrm{input}}},\\ G&=\mathrm{Adapter}(Z)\;\in\;\mathbb{R}^{B\times T^{\prime}\times D_{\mathrm{adapter}}},\\ L_{\mathrm{lip}}&=\mathrm{CTCHead}(G)\;\in\;\mathbb{R}^{B\times T^{\prime}\times N_{\mathrm{phonemes}}}\end{split}$ (7)
This processes the same video to produce complementary speech-based token predictions. These outputs consist of all phonemes in the English language.
Multimodal Feature Fusion
To integrate information from manual (signs, fingerspelling) and non-manual (lips) modalities while addressing their temporal asynchronicity, we employ a transformer-based fusion module.
First, we project the token-level embeddings for each modality into a shared fusion dimension $D$:
$L^{\prime}_{\mathrm{sign}}$ $=L_{\mathrm{sign}}W_{\mathrm{sign}}+b_{\mathrm{sign}},$ (8)
$L^{\prime}_{\mathrm{fs}}$ $=L_{\mathrm{fs}}W_{\mathrm{fs}}+b_{\mathrm{fs}},$
$L^{\prime}_{\mathrm{lip}}$ $=L_{\mathrm{lip}}W_{\mathrm{lip}}+b_{\mathrm{lip}}$
Next, we create a unified manual representation, $M$, by combining the projected manual features. This combination is weighted by Gumbel-Softmax scores ($g_{\mathrm{sign}},g_{\mathrm{fs}}$) from a sequence-type classifier that determines the active modality for each token. A learned null vector, $\tilde{n}\in\mathbb{R}^{D}$, handles rest poses, weighted by $g_{\mathrm{rest}}$:
$M=g_{\mathrm{sign}}L^{\prime}_{\mathrm{sign}}+g_{\mathrm{fs}}L^{\prime}_{\mathrm{fs}}+g_{\mathrm{rest}}\tilde{\mathbf{n}}\quad\in\mathbb{R}^{B\times T\times D}$ (9)
A gating mechanism then fuses this unified manual representation $M$ with the non-manual lip features $L^{\prime}$. An element-wise fusion gate $G$ is computed as:
$G=\sigma\bigl{(}MW_{g}+b_{g}\bigr{)}\quad\in\mathbb{R}^{B\times T\times D}$ (10)
where $\sigma$ is the sigmoid function. The final fused representation, $H_{\mathrm{fused}}$, is produced by:
$H_{\mathrm{fused}}=G\odot M+(1-G)\odot L^{\prime}_{\mathrm{lip}}$ (11)
where $\odot$ denotes element-wise multiplication.
Finally, $H_{\mathrm{fused}}$ is passed through a standard $L$-layer transformer encoder to model temporal dependencies. The process is defined as:
$H^{(0)}$ $=H_{\mathrm{fused}}W_{p}+b_{p},$ (12)
$H^{\prime(\ell)}$ $=\mathrm{LayerNorm}\left(H^{(\ell-1)}+\mathrm{MHSA}\left(H^{(\ell-1)}\right)\right),$
$H^{(\ell)}$ $=\mathrm{LayerNorm}\left(H^{\prime(\ell)}+\mathrm{FFN}\left(H^{\prime(\ell)}\right)\right),$
$Z$ $=H^{(L)}W_{o}+b_{o}$
where $\ell=1,\dots,L$. The output $Z\in\mathbb{R}^{B\times T\times D}$ is the final sequence of contextualised embeddings.
Model Modality Test on How2Sign
Pose RGB BLEU-1 $\uparrow$ BLEU-2 $\uparrow$ BLEU-3 $\uparrow$ BLEU-4 $\uparrow$ ROUGE $\uparrow$
GloFE-VN (Lin et al. 2023) 14.9 7.3 3.9 2.2 12.6
MSLU (Zhou et al. 2024) 20.1 7.7 - 2.4 17.2
SLT-IV (Tarrés et al. 2023) 34.0 19.3 12.2 8.0 -
C2RL (Chen et al. 2024a) 29.1 18.6 12.9 9.4 27.0
FLa-LLM (Chen et al. 2024b) 29.8 19.0 13.3 9.7 -
YouTube-ASL (Uthus, Tanzer, and Georg 2023) 37.8 24.1 16.9 12.4 -
SignMusketeers (Gueuwou et al. 2025) 41.5 27.2 19.3 14.3 -
Uni-Sign (Li et al. 2025) 40.2 27.1 19.7 14.9 36.0
SSVP-SLT (Rust et al. 2024) 43.2 28.8 20.8 15.5 38.4
MultiStream-LLM (Ours) 46.0 33.5 26.8 23.5 41.2
Table 1: Modality (pose / RGB) and n-gram BLEU / ROUGE comparison on How2Sign. MultiStream-LLM achieves the best scores across all metrics.
Training and Optimization
Rather than traditional end‐to‐end training, we employ a staged training schedule in which each component is optimized separately. Firstly, the Sequence Type Classifier and all three Modality-Specific Predictors are pre-trained. Then all the pre-trained stages are recombined and frozen to train the Fusion Transformer. Finally, the LLM is fine-tuned to recreate sentences from the Fusion Transformer output tokens. All stages use the Adam optimizer (Kingma and Ba 2014).
  1. 1.
    Sequence Type Classifier
    Train the ViT–based classifier with standard cross‐entropy:
    $\mathcal{L}_{\mathrm{type}}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{c=1}^{N_{\mathrm{classes}}}y_{i,c}\,\log\hat{y}_{i,c}$ (13)
    where $\hat{y}_{i,c}$ is the softmax score for class $c$ on example $i$.
  2. 2.
    Modality-Specific Predictors
    Separately train the fingerspelling, sign-language, and lip-reading branches using CTC loss:
    $\mathcal{L}_{m}=\mathrm{CTC}\bigl{(}L_{m},T_{m}\bigr{)},\quad m\in\{\mathrm{fs},\mathrm{sign},\mathrm{lip}\}$ (14)
    where $L_{m}$ are the per‐frame logits from branch $m$ and $T_{m}$ the corresponding ground-truth token sequence.
  3. 3.
    Fusion Transformer
    With the predictors frozen, train the fusion transformer and CTC head on the fused logits $\hat{Y}_{\mathrm{fusion}}$:
    $\mathcal{L}_{\mathrm{fusion}}=\mathrm{CTC}\bigl{(}\hat{Y}_{\mathrm{fusion}},T_{\mathrm{gt}}\bigr{)}$ (15)
    where $T_{\mathrm{gt}}$ is the target token sequence for the full translation.
  4. 4.
    LLM Fine-Tuning
    If paired text is available, fine-tune the prefix‐tuning or full LLM with cross-entropy:
    $\mathcal{L}_{\mathrm{LLM}}=-\frac{1}{N}\sum_{t=1}^{N}z^{*}_{t}\,\log\hat{z}_{t}$ (16)
    where $\hat{z}_{t}$ is the LLM’s output distribution at step $t$ and $z^{*}_{t}$ the one-hot ground-truth token.
The total loss can be calculated as follows:
$\mathcal{L}_{\mathrm{total}}$ $=\underbrace{\mathcal{L}_{\mathrm{type}}}_{\text{Type Classification}}+\sum_{m\in\{\mathrm{fs},\mathrm{sign},\mathrm{lip}\}}\underbrace{\mathcal{L}_{m}}_{\text{Predictor }m}$ (17)
$\quad+\underbrace{\mathcal{L}_{\mathrm{fusion}}}_{\text{Fusion}}+\underbrace{\mathcal{L}_{\mathrm{LLM}}}_{\text{LLM Fine-Tuning}}$
Experiments
This section discusses the datasets used for training and evaluation, and the metrics employed to assess model performance.
Datasets
We pre-trained our model using the Youtube-ASL dataset and ChicagoFSWild+, then fine-tuned it on How2Sign. We then evaluated our proposed framework on How2Sign and ChicagoFSWild+.
YouTube-ASL (Uthus, Tanzer, and Georg 2023):
A large-scale ASL dataset comprising 60K videos and 1,000 hours sourced from YouTube, annotated with sentence-level transcriptions. This dataset provides diverse signing conditions, including variations in background, lighting, and signer appearance. The dataset features over 2,500 unique signers, ensuring that the model generalizes well across different users.
ChicagoFSWildPlus (B. Shi and Livescu 2019):
A dataset containing 55,232 ASL finger-spelling sequences performed by 260 signers containing in-the-wild videos of fingerspelling sequences. It includes multiple signers, varied environments, and occlusions, which makes it ideal for evaluating the robustness of fingerspelling detection.
How2Sign (Duarte et al. 2021):
How2Sign is a large‐scale, continuous ASL corpus derived from 2,456 instructional “How2” videos, totaling over 8080 hours of footage. Recordings were made in two settings: a Green-Screen studio (79.1 h over 2,529 videos) and the Panoptic studio (2.96 h over 124 videos). Eleven signers (5 hearing ASL interpreters, 2 hard-of-hearing, 4 Deaf) produced 35,19135,191 sentence‐level clips (average 162 frames/5.4 s, 17 words), yielding a vocabulary of over 16,000 English words.
Evaluation Metrics
We evaluated model performance using the following metrics:
Letter Accuracy (Garofolo et al. 1993):
Evaluates performance at the character level, particularly important for fingerspelling. It captures errors in individual letter predictions.
Sequence Accuracy:
The percentage of correctly predicted sequences. This metric reflects the model’s ability to classify the type of sequence, be it Sign, Fingerspelling or Nothing.
BLEU Score (Papineni et al. 2002):
Assesses the quality of generated sequences by comparing them to reference translations. A higher BLEU score indicates better alignment with human-generated annotations.
ROUGE (Ganesan 2018):
Measures the overlap of n-grams, longest common subsequences, and word sequences between a generated summary (or translation) and one or more reference texts, prioritizing recall of relevant content.
Model Letter Accuracy
FS Recognition (Shi et al. 2019b) 41.2%
Iterative Attention (Shi et al. 2019a) 46.7%
Pannattee et al. (Pannattee et al. 2021) 48.0%
Gajurel et al. (Gajurel, Zhong, and Wang 2021) 48.36%
TDC-SL (Papadimitriou and Potamianos 2020) 50.0%
FS Attention (Kabade et al. 2023) 57.84%
FSS-Net (Shi et al. 2022a) 64.4%
CtoML (Li et al. 2023) 54.9%
FS PoseNet (Fayyazsanavi, Nejatishahidin, and Kosecka 2023) 71.1%
MultiStream-LLM (Ours) 73.2%
Table 2: Performance comparison of Letter Accuracy on ChicagoFSWildPlus; our approach achieves the highest score of 73.2%.
Results
In this section we compare the performance of MultiStream-LLM with several state-of-the-art models from recent literature, we achieved a sequence classification accuracy of 99.2%. We split the results into two tables, Table 1 shows the comparison between our approach and other approaches in SLT given in BLEU-1 through BLEU-4 and Table 2 shows the comparison for fingerspelling in Letter Accuracy (%).
Ablation Study
In this section, we discuss the different ablation studies performed to demonstrate the contribution of the various components in the architecture.
Model Variants:
The first ablation was to understand the contribution of each modality and component, we conducted an ablation study by selectively disabling key features of MultiStream-LLM. Table 3 presents the results.
Model Variant Letter Acc. (%) BLEU-4
Without Lipreading 54.3 16.6
Without LLM 58.8 7.8
Full Model 73.2 23.5
Table 3: Comparison of Letter Accuracy and BLEU-4 score when different components are removed from the model. The table shows that removing the LLM drops BLEU-4 down to 7.8 and the Letter Accuracy to 58.9%. When the Lipreading is removed the letter accuracy drops down to 54.3% and the BLEU-4 down to 16.6.
The ablation results highlight the importance of each component. Removal of the lip-reading model leads to a decrease in Letter Accuracy (18.9%) and a decrease in the BLEU score (6.9), indicating that lip cues play a crucial role in disambiguating finger-spelled words and an important role in clarifying Sign Language. Removing the LLM and leaving the CTC head results in the largest performance drop, 14.4% for Letter Accuracy and 15.7 for BLEU-4, showing the importance the LLM has in sentence reconstruction.
Varying LLMs:
The next ablation was to compare the effectiveness of different LLMs, fully fine-tuned with different parameter sizes. Table 4 shows the results of this study.
Model Parameters (B) Letter Acc. (%) BLEU-4
Llama 3.2-1B 1.0 60.9 12.1
Llama 3.2-3B 3.0 64.8 16.4
Llama 2-7B 7.0 73.2 23.5
Table 4: Comparison of Letter Accuracy and BLEU-4 scores for different LLMs (Parameters in billions) tested with our model. The model with the worst performance is the 1 Billion parameter model with a Letter Accuracy of 60.9% and BLEU of 12.1. The 7 Billion parameter model achieves the best results with a Letter Accuracy of 73.2% and BLEU of 23.5.
Glossing Methods:
This ablation compares different forms of pseudo-glossing to determine their effects on model performance, these include:
  • Random Glossing: N-number of words are randomly selected from the list of possible labels such that every label contains at least n of these Glosses
  • Probability Glossing: N-number of words are selected from the list of possible labels based on the frequency they show up (if two entries have the same frequency then the first label alphabetically is selected)
  • Pseudo-Glossing: N-number of words are selected based on their importance to sentence structure
    • LLM Glossing: An LLM prepared with knowledge of ASL sentence structure selects the most important words from the list of available words
    • Keyword-Glossing: Traditional Pseudo glossing, where nouns, adjectives, vers, etc are selected using Wordnet.
Glossing Method BLEU-4
Baseline
No Glossing 9.6
Random Glossing
2 Words 1.3
3 Words 6.7
5 Words 15.2
7 Words 16.0
Probability Glossing
1000 words 5.6
2000 words 14.8
3000 words 20.1
5000 words 17.6
Pseudo-Glossing
Keyword Glossing 22.4
LLM Glossing 23.5
Table 5: Comparison of different glossing methods and their effects on BLEU-4 score. The table shows that the optimum number of Random glosses per label is 5, the optimum number of Probability based Glosses is 3000. We also see that the worst option is 2 random glosses per label
The results of Table 5 show that the optimum method of glossing is LLM glossing, where an LLM determines which tokens are the most crucial for sentence reconstruction. For example: the LLM dynamically deciding that in one sentence an auxiliary verb is crucial, whereas in another it might be safely dropped. Where as a static approach such as using WordNet is unable to do that. We can also see that not using a Glossing method still produces superior results to the limited Glossing methods. For example: Randomly Glossing 2 words in the label “I’m going to teach you how to make one of my favourite paper airplanes.” would give insufficient information for the LLM to recreate the sentence.
Asynchronous Fusion:
The final ablation focuses on the effectiveness of the fusion model. The aim is to quantify how sensitive our fusion mechanism is to the temporal mis‐alignment that naturally occurs between hand movements (manual signals) and mouth movements (non‐manual cues) in real sign language. To quantify this we test with four different temporal shifts:
  1. 1.
    No shift: Directly fuse frame $t$ of lips with frame $t$ of hands, no temporal shift at all
  2. 2.
    $\Delta\pm$ 5 frames: fuse frame t of hands with lip $t\pm 3$ frames
  3. 3.
    $\Delta\pm$ 10 frames: fuse frame t of hands with lip $t\pm 5$ frames
  4. 4.
    Learned Alignment: The model learn an optimal per‐time fusion gating
Variant Letter Acc. (%) BLEU-4
No shift 68.3 17.2
$\Delta+5$ 62.8 12.4
$\Delta-5$ 70.6 18.1
$\Delta+10$ 55.7 7.8
$\Delta-10$ 64.9 15.6
Learned alignment 73.2 23.5
Table 6: Comparison of different temporal alignments between manual and non-manual streams and their effect on Letter Accuracy and BLEU-4 score. The table shows that the Learned Alignment provides the best scores and by delaying the fusion by 10 frames drops the score the most.
Conclusion
In this paper, we introduced MultiStream-LLM, a novel framework for gloss-free sign language translation that integrates lip-reading, fingerspelling detection, and continuous sign language recognition into a unified pipeline. By leveraging a transformer-based architecture and an LLM-driven sequence prediction head, MultiStream-LLM demonstrated significant improvements over conventional CTC-based approaches. Our model effectively addresses the challenges posed by asynchronous modalities and enhances translation accuracy through multimodal fusion.
Supplementary Material
Translation Examples
Here we show translations using our model on the How2Sign dataset. Example translations can be seen in Table LABEL:Table:Translation, where we evaluate the outputs from each of our “expert” predictors: continous sign, fingerspelling, and lip reading as well as the fused final sentence. From the first () examples we can see that our model is capable of learning the difference between continuos sign language and fingerspelling; and that the fusion model and LLM successfully captures semantic information in sign language, generating sentences that accurately map to the reference labels. However, we also observed occasional failure cases, for example 5 where the model failed to correctly generate complex non-English names; and example 6 where the model generated linguisticly similar but different translations.
Table 7: Qualitative examples of our multi-modal translation system. Our method effectively fuses continuous signing, isolated fingerspelling, and lip-read phonemes. Correctly generated words are highlighted in green and incorrectly generated words are highlighted in red
Modality / Component Output / Transcription
Input: Sign My name [NAME] and let get down to it.
Input: Fingerspelling D E A N - H A L E
Input: Lip Reading (Phonemes) /M/ /AY/ /N/ /EY/ /M/ /IH/ /Z/ /D/ /IY/ /N/ /HH/ /EY/ /L/ /AH/ /N/ /D/ /L/ /EH/ /T/ /S/ /G/ /EH/ /T/ /D/ /AW/ /N/ /T/ /UW/ /IH/ /T/
Our Method (Combined) My name is Dean Hale and let’s get down to it.
Reference My name is Dean Hale let get down it.
Input: Sign We have [NAME] standing about hip width apart.
Input: Fingerspelling E M I L Y
Input: Lip Reading (Phonemes) /W/ /IY/ /L/ /HH/ /AE/ /V/ /EH/ /M/ /AH/ /L/ /IY/ /S/ /T/ /AE/ /N/ /D/ /IH/ /NG/ /AH/ /B/ /AW/ /T/ /HH/ /IH/ /P/ /W/ /IH/ /D/ /TH/ /AH/ /P/ /AA/ /R/ /T/
Our Method (Combined) We’ll have Emily standing about hip width apart.
Reference We’ll have Emily standing about hip width apart.
Input: Sign In Chicago [NAME]
Input: Fingerspelling F R E D - A B - B O B - F O O S E
Input: Lip Reading (Phonemes) /IH/ /N/ /SH/ /AH/ /K/ /AA/ /G/ /OW/ /B/ /AY/ /F/ /R/ /EH/ /D/ /AE/ /B/ /AH/ /N/ /D/ /B/ /AA/ /B/ /F/ /AA/ /S/
Our Method (Combined) In Chicago by Fred Ab and Bob Fosse.
Reference In Chicago by Fred Ab and Bob Fosse.
Input: Sign Welcome ladies gentlemen you about see story murder, greed, corruption, exploitation, treachery adultery
Input: Fingerspelling -
Input: Lip Reading (Phonemes) /W/ /EH/ /L/ /K/ /AH/ /M/ /L/ /EY/ /D/ /IY/ /Z/ /JH/ /EH/ /N/ /N/ /T/ /AH/ /L/ /M/ /IH/ /N/ /AH/ /B/ /AW/ /T/ /T/ /UW/ /S/ /IY/ /S/ /T/ /AO/ /R/ /IY/ /AH/ /V/ /M/ /ER/ /D/ /ER/ /G/ /IY/ /D/ /K/ /ER/ /P/ /SH/ /AH/ /N/ /EH/ /K/ /S/ /P/ /L/ /OY/ /T/ /EY/ /SH/ /AH/ /N/ /T/ /R/ /EH/ /CH/ /ER/ /IY/ /AH/ /N/ /D/ /AH/ /D/ /AH/ /L/ /T/ /IY/
Our Method (Combined) Welcome ladies and gentlemen you’re about to see a story of murder, greed, corruption, exploitation, treachery and adultery.
Reference Welcome ladies and gentlemen you’re about to see a story of murder, greed, corruption, exploitation, treachery and adultery.
Input: Sign Thank you everybody [NAME].
Input: Fingerspelling D A V I D O D O N N E L L
Input: Lip Reading (Phonemes) /TH/ /AE/ /NG/ /K/ /Y/ /UW/ /EH/ /V/ /R/ /IY/ /B/ /AA/ /D/ /IY/ /IH/ /M/ /D/ /EY/ /V/ /IH/ /D/ /OW/ /D/ /AA/ /N/ /AH/ /L/
Our Method (Combined) Thank you everybody, I’m David Donnell
Reference Thank you everybody, I’m David O’Donnell.
Input: Sign you see how filling in brows really frames face, lifts eye gives more polished look.
Input: Fingerspelling -
Input: Lip Reading (Phonemes) /S/ /OW/ /Y/ /UW/ /K/ /AE/ /N/ /S/ /IY/ /HH/ /AW/ /F/ /IH/ /L/ /IH/ /NG/ /IH/ /N/ /DH/ /AH/ /B/ /R/ /AW/ /Z/ /R/ /IH/ /L/ /IY/ /F/ /R/ /EY/ /M/ /Z/ /DH/ /AH/ /AY/ /AH/ /N/ /D/ /G/ /IH/ /V/ /Z/ /AH/ /M/ /AH/ /CH/ /M/ /AO/ /R/ /P/ /L/ /IH/ /SH/ /T/ /L/ /UH/ /K/
Our Method (Combined) So you can see how filling in the eyebrows frames the face, lifts the eye and gives a much more clean look
Reference So you can see how filling in the brows really frames the face, lifts the eye and gives a much more polished look.
Translation Visualizations
This section presents visualizations of our model’s translation process on the How2Sign dataset. At the top of each figure, we display the ground truth reference sentence. Below that, we include representative video frames extracted from the input sequence.
The central graph shows the model’s modality branch activation probabilities over time: blue indicates continuous signing, red corresponds to fingerspelling, and green represents a rest period. These activations reflect how the model dynamically routes segments through different specialized pathways during inference.
At the bottom of each figure, we show the translated output corresponding to each temporal segment.
In Figure 2, we present a detailed visualization of the translation process for the sentence: “Hi, my name is Robert Segundo and today I’m going to teach you how to make one of my favorite paper airplanes, the simple one.” This example illustrates how our model dynamically integrates information from multiple modalities over time.
As the video progresses, the model identifies fingerspelled segments (e.g., “Robert Segundo”) with high confidence, while relying on the continuous sign branch for fluent phrases such as “Hi, my name is…” and “the simple one.” The change in class probabilities reflects the shifting visual cues within the video and demonstrates the model’s ability to route information to the appropriate stream. At the bottom of the figure, we show the corresponding translation for each segment, which aligns closely with the reference, indicating that the model correctly interprets both content and context.
In Figure 3, we visualize the translation of the sentence: “Hello, my name is Eric White and today we’re going to talk about how to edit the text in your MySpace page.” This example further highlights the model’s ability to track modality relevance across time. We display class probabilities, which reveal transitions from fingerspelling (for “Eric White”) to general signing and instructional language. The model effectively attends to fingerspelled inputs and transitions smoothly into broader sentence construction, maintaining accuracy throughout. The final translation remains semantically faithful to the reference, illustrating the robustness of our multi-stream approach even in longer, information-dense utterances.
Figure 2: Example of the model’s sequence classification and sentence outputs from “Hi, my name is Robert Segundo and today I’m going to teach you how to make one of my favorite paper airplanes, the simple one.” in the How2Sign dataset (Duarte et al. 2021). The graph illustrates the model’s class probability distribution for different sequences from visual input, alongside the final text predictions. The model is able to identify all instances of fingerspelling, signing and rest; in this example the entire sentence is predicted correctly.
Figure 3: Example of the model’s sequence classification and sentence outputs from “Hello, my name is Eric White and today we’re going to talk about how to edit the text in your My Space page.” in the How2Sign dataset (Duarte et al. 2021). The graph illustrates the model’s class probability distribution for different sequences from visual input, alongside the final text predictions. The model is able to identify both instances of fingerspelling and all instances of continuos signing; in this example the entire sentence is predicted correctly.